The Alpine House, Inc., is a large retailer of snow skis. The company assembled the information shown below for the quarter ended March 31: Amount Sales $ 1,435,000 Selling price per pair of skis $ 410 Variable selling expense per pair of skis $ 45 Variable administrative expense per pair of skis $ 17 Total fixed selling expense $ 150,000 Total fixed administrative expense $ 100,000 Beginning merchandise inventory $ 75,000 Ending merchandise inventory $ 105,000 Merchandise purchases $ 295,000 Required: 1. Prepare a traditional income statement for the quarter ended March 31. 2. Prepare a contribution format income statement for the quarter ended March 31. 3. What was the contribution margin per unit?
Answer #1

Answer:

1 traditional income statement for the quarter ended March 31.
Amount in $
Sales                   14,35,000
Less: Cost of goods Sold                     2,65,000
gross margin                   11,70,000
less: Selling & Admin Expenses
Selling Expenses             3,07,500
Administrative Expenses             1,59,500
                    4,67,000
Net operating Income                     7,03,000
Cost of goods sold = Opening merchandise + Purchases - Closing Merchandise =$75000+$295000-$105,000 =$265,000
Total pairs of skis = Total sales / Per pair skis price =$1435000/$410 =3500
Selling Expenses =$45*3500+ $150,000 =307500
Administrative Expenses =$17*3500 +$100000 =$159,500
2 contribution format income statement for the quarter ended March 31.
Sales                   14,35,000
less: Variable Expenses
Cost of goods sold             2,65,000
Selling expenses             1,57,500
Administrative Expenses                59,500
                    4,82,000
Contribution margin                     9,53,000
Less: Fixed Expenses
Selling expenses             1,50,000
Administrative Expenses             1,00,000
                    2,50,000
Net operating income                     7,03,000
Cost of goods sold = Opening merchandise + Purchases - Closing Merchandise =$75000+$295000-$105,000 =$265,000
Selling Expenses =$45*3500=$157500
Administrative Expenses =$17*3500 =$63000
3 the contribution margin per unit?
Total contribution margin $953000
Total number of pairs                  3,500
Contribution margin per pair ($953,000/3500)=$272.29